How many quarts of pure antifreeze must be added to 6 quarts of a 10% solution to obtain a 20% antifreeze soltuion? (Round to the nearest tenth.)
Answer by mananth(16946)

How many quarts of pure antifreeze must be added to 6 quarts of a 10% solution to obtain a 20% antifreeze soltuion? (Round to the nearest tenth.)
Let the quantity of pure antifreeze added be x
pure(100%)--------------- 10%------------------------- 20 % mix
x-------------------------6-------------------------- x+6 quarts
1x+0.6x=0.2(x+6)
1.6x= 0.2x + 1.2
1.6x-0.2x= 1.2
1.4x= 1.2
x= 1.2/1.4
0.85 quarts